Circular No. 8901

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 W3 (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation tabulated below), and posted on the Minor Planet Center’s NEOCP webpage, has been found to show cometary appearance by several astrometric CCD observers, including R. Holmes (Charleston, IL, U.S.A., 0.61-m f/4.0 astrograph; Dec. 1.31 UT; measurer S. Foglia; five stacked images show a coma diameter of 6′′ and no tail); Q.-z. Ye (Department of Atmosphere Science, Sun Yat-sen University, Guangzhou, China; Lulin Sky Survey 0.41-m Ritchey-Chrétien reflector; Dec. 2.6; four combined 60-s unfiltered exposures show a small bright core < 2′′ in diameter with a total coma diameter of ∼ 7′′–8′′ that was slightly elongated in p.a. 90^{o}); J. Young (Table Mountain, 0.61-m f/16 Cassegrain reflector; Dec. 3.3; very round 6′′ coma with bright central condensation and a possible hint of a broad, stubby 10′′ tail in p.a. 190^{o}); and E. Guido and G. Sostero (Castellammare di Stabia, Italy, remotely using a 0.25-m f/3.4 reflector located near Mayhill, NM, U.S.A.; Dec. 3.5; 60 co-added unfiltered exposures show a coma nearly 8′′ in diameter, slightly elongated toward the northeast).

2007 UT  α_{2000}  δ_{2000}  Mag.
Nov. 29.31590  6^{h}09^{m}41.^{s}63  +72^{o}37′01′′.9  19.6

The available astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2007-X13.

T = 2008 June 2.793 TT  ω = 112.^{o}585
Ω = 73.074 }2000.0
q = 1.77775 AU  i = 78.679

V598 PUPPIS R. W. Russell, R. J. Rudy, and D. K. Lynch, The Aerospace Corpora-

tion; and C. E. Woodward, University of Minnesota, report 0.8- to 2.5-μm spectroscopy of this nova (cf. IAUC 8898, 8899) using the Infrared Telescope Facility (+ SpeX) on Nov. 30.52 UT. V598 Pup was in its coronal stage with no low excitation lines present. [Si VI] 1.96-μm and [Ca VIII] 2.32-μm were strong, along with He I 1.0830-μm and the H I lines. At least two of the unidentified novae lines were present (1.55 and 2.09 μm). The FWHM of the lines was ≈ 2400 km/s, and there was no evidence of thermal emission from dust.

2007 December 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green